OpenStudy (annie96):

To get from the post office to the library, Sindy walked due north and then due east. The distance Sindy walked east is 20 meters more than the distance she walked north. If Sindy were able to walk directly from the post office to the library in a straight line, she would have walked 150 meters. Suppose x represents the distance Sindy walked north. Which equation can be used to find x? x^2+(x+20)^2=150^2 202=x^2-150^2 x+(x+20)=150 x^2+(x+150)^2=20^2

OpenStudy (annie96):

is it a?

OpenStudy (mathmale):

Although some people can do this problem without sketching the situation described, it's easier to visualize the distances involved if you do sketch it. Are you willing to do that? Three different distances (lines) are mentioned in this problem statement. They form which geometric figure? Which well-known theorem would you apply to the given info to solve this problem?

OpenStudy (annie96):

yes

OpenStudy (mathmale):

OK. Please sketch this situation. Use the Draw utility (click on button, below). Label everything.

OpenStudy (annie96):

x = north x + 20 = east A triangle with north and east as side lengths and 150 as hypotenuse length. north^2 + east^2 = 150^2 x^2 + (x + 20)^2 = 150^2